OpenStudy (anonymous):

What are the solutions of the quadratic equation below? -7x^2 - 23x + 10 = 0

OpenStudy (anonymous):

so which ones are a,b, and c

OpenStudy (shamil98):

ax^2 + bx + c = 0

OpenStudy (shamil98):

x = -b +- √b^2 - 4ac / 2a

OpenStudy (anonymous):

First rewrite the equation as 7x^2 + 23x - 10 = 0 Then use the quadratic formula with a = 7, b = 23, c = -10

OpenStudy (jhannybean):

\[\large -7x^2 - 23x + 10 = 0\]\[\large x^2 -23x -70=0\]\[\large (x^2 -23x)-70=0\]\[\large (x^2-23x+\frac{529}{4}) -70-\frac{529}{4}=0\]\[\large (x-\frac{23}{2} )^2 -\frac{809}{4}=0\]\[\large \sqrt{(x-\frac{23}{2})^2} = \pm\sqrt{\frac{809}{4}}\]\[\large x-\frac{23}{2} =\pm \frac12 \sqrt{809}\]\[\large x= \frac{23}{2} \pm \frac12 \sqrt{809}\]

OpenStudy (anonymous):

does something have to be reduced?? @Jhannybean

OpenStudy (anonymous):

and i think something has to be negative

OpenStudy (jhannybean):

Yes, you will get two values. \[\large x= 23 + \frac{1}{2}\sqrt{809}\]\[\large x= 23-\frac{1}{2}\sqrt{809}\]

OpenStudy (jhannybean):

http://www.wolframalpha.com/input/?i=x%5E2+-23x+-70

OpenStudy (anonymous):

none of those match the options.. A. 23+- √809 / -14 B. 23+-√809 / -7 C. -23+-√809 / -14 D. -23+-√809 / -7

OpenStudy (jhannybean):

the quadratic formula \[\large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]\[\large -7x^2 - 23x + 10 = 0\]\[\large a=-7 \ , \ b=-23 \ , \ c =10\]Input this into the formula. \[\large x=\frac{-(-23)\pm \sqrt{(-23)^2 -4(-7)(10)}}{2(-7)}\]\[\large x=\frac{23 \pm \sqrt{529 +280}}{-14}\]\[\large x=\frac{23 \pm \sqrt{809}}{-14} \]\[\large x= \frac{23}{14} \pm \sqrt{809}\]

OpenStudy (jhannybean):

In the method i had posted the first time, i forgot to redivide everything by 7, since we had multiplied it over. The only thing i wasmissing. Sorry about that.

OpenStudy (anonymous):

so it's A right!

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

you have no clue how much this means to me that you showed me how to do this! thank you!

OpenStudy (jhannybean):

No problem :)